OpenStudy (anonymous):

integral 23/xln3x dx

OpenStudy (kirbykirby):

\[\int \frac{23}{x}ln(3x)dx\]?

OpenStudy (kirbykirby):

is that the integral, just to be sure

OpenStudy (anonymous):

Substitute \(u=\ln(3x)~\Rightarrow~du=\dfrac{3}{3x}~dx=\dfrac{1}{x}~dx\). The integral changes to \[23\int u~du\]